OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? answers: 430 720 1150 1300

OpenStudy (texaschic101):

If your curious and want to learn, this is how you set it up... o = one step equations t = two step equations o + t = 1150 -->t = 1150 - o 3o - 2t = 1300 Now you would sub 1150 - o in for t in the 2nd equation and then solve for o.

OpenStudy (anonymous):

thank you so much!